OpenStudy (anonymous):

Let f(x) be odd and integral between -4 to 8 f(x)dx=1. Integral Between 4 to 8 f(x)dx=?

OpenStudy (anonymous):

second time i have seen this problem and i couldn't get it the first time seems like it could be anything

OpenStudy (anonymous):

I know that there's a specific way to do it, but problem is i don't know how to apply it. xD

OpenStudy (anonymous):

ooh doh

OpenStudy (anonymous):

\[\int_{-4}^4f(x)dx=0\] because the function is odd!!

OpenStudy (anonymous):

Apparently It's not the right answer. This is what I got a hint for.

OpenStudy (anonymous):

jeez sometimes i think my brain went on vacation

OpenStudy (anonymous):

yea so the part from -4 to 4 contributed nothing

OpenStudy (anonymous):

the part from 4 to 8 is one

OpenStudy (anonymous):

Sometimes I want to cry in a corner.

OpenStudy (anonymous):

lol me too

OpenStudy (anonymous):

\[\int_{-4}^8=\int_{-4}^4+\int_4^8=0+\int_4^8=1\]

OpenStudy (anonymous):

That seems right. lol actually yeah!

OpenStudy (anonymous):

oh yeah, it is right

OpenStudy (anonymous):

i would bet \(\$87\) on it

OpenStudy (anonymous):

Hahaha Thanks :)

myininaya (myininaya):

And if f is odd then F is even so F(4)=F(-4) \[\int\limits_{-4}^{8}f(x) dx=F(8)-F(-4)=F(8)-F(4)=1\]

myininaya (myininaya):

\[\text{ so } \int\limits_{4}^{8}f(x) dx=1 \]